OpenStudy (anonymous):

what is the sum of the arithmetic sequence 17,13,9... if there are 40 terms ?

OpenStudy (amistre64):

we can either find the 40th term or use the set up for the 40th term; either way is suitable

OpenStudy (amistre64):

an = a1 + d(n-1) where a1 is the first term; d is the difference between terms; and n is the nth term we want

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

so help me out; our first term (a1) is?

OpenStudy (anonymous):

17?

OpenStudy (amistre64):

yes and since we want the nth term to be 40 lets go ahead and use that as well a40 = 17 + d(40-1) all we need now is the "d" 17 + d = 13 should suffice

OpenStudy (anonymous):

d= 4

OpenStudy (amistre64):

well, -4 :) 17 + (-4) = 13 ... sooo a40 = 17-4(39) is what we need to use in the summation formula that i presented last time

OpenStudy (amistre64):

a40 = -139 if my calculator hasnt lied to me so lets fill in the summation with these: \[S=\frac{n(a_1+a_{40})}{2}\] \[S=\frac{40(17-139)}{2}\]

OpenStudy (anonymous):

-2440

OpenStudy (amistre64):

-2440, i agree

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

youre welcome